# Flock: A Knowledge Graph Foundation Model via Learning on Random Walks

### v1

@@ 4.1 Flock
4.2 Theoretical properties of Flock
Expressivity. Following the notion of probabilistic expressivity introduced by Abboud et al. (2021), we say that a Flock model $X_{\theta}$ is a universal approximator of link invariant functions over $\mathbb{K}_{n,m}$ if for any link invariant $\varphi:\mathbb{K}_{n,m}\rightarrow(V\times R\times V\rightarrow[0,1])$ and any $\epsilon,\delta>0$, there exists a choice of the network parameters $\theta$ and the length of the sampled random walks $l$, such that:
@@ 4.2 Theoretical properties of Flock
Any Flock model with components as outlined in this section, and detailed in Appendix B is invariant in probability.
@@ 5 Experiments
We evaluate Flock over a wide range of KGs with both entity and relation prediction tasks, aiming to answer the following questions:
@@
    How does performance scale with the pretraining graph mix and the size of test-time ensemble?
@@
    What is the impact of test-time adaptation of walk counts?
  • Q5.
    How do current KGFMs perform under noise injection, and how do they compare with Flock? (see Appendix F)
In addition to these experiments, we report the computational complexity of Flock in Appendix D and conduct a detailed scalability analysis in Appendix E. We report ensembled predictions by averaging $P$ independent stochastic forward passes to reduce variance. Please refer to the detailed per-dataset choices in Table 22.
5.1 Synthetic datasets experiments
@@
Figure 3: Example KG from Petals. KGFMs with relational invariants must equate blue $r_{1}$ and red $r_{2}$, thus predicting the same scores for both dashed queries with $r_{0}$.
Setup. We construct a synthetic benchmark Petals to validate the limitations of KGFMs following the node-relation equivariance (Q1). The Petals benchmark contains $220$ instances, with each instance including the following elements: 1) a KG $G=(V,E,R)$, which consists of a ‘central’ node $s\in V$, a ‘stem’ $T\subset V$ with query relation $r_{0}\in R$, and multiple cyclic ‘petals’, each ‘colored’ with a different pair of relations in $R\setminus\{r_{0}\}$, 2) an entity prediction query $(h,r_{0},?)$ with $h\in\{s\}\cup T$, and 3) two candidate targets $t_{1}$ and $t_{2}$ from the same ‘petal’, located at the same distance from $s$. An example is shown in Figure 3. See Appendix A for more details.
Petals is designed such that each instance always admits non-trivial automorphisms, meaning that swapping relations occurring in the same ‘petal’ will result in an isomorphic KG. Consequently, any model computing relation invariants will not be able to distinguish between potential links $(s,r_{0},t_{1})$ and $(s,r_{0},t_{2})$. However, the samples are constructed so that these links are not isomorphic from the graph perspective, making them distinguishable for general link-invariant functions. We say a model successfully solves an instance if it can classify $(s,r_{0},t_{1})$ as true and $(s,r_{0},t_{2})$ as false, and we measure the accuracy in the experiment.
Table 1: Accuracy of KGFMs of our Petals benchmark.
@@
We trained Ultra (Galkin et al., 2024), $\textsc{Motif}({\mathcal{F}}_{\text{Path}}^{3})$ (Huang et al., 2025), Trix (Zhang et al., 2024), and Flock from scratch and validated them on the training instances.
Results. We present the results in Table 1. As expected, all existing KGFMs relying on learning deterministic relational invariants fail to distinguish between the candidate target triplets completely, achieving $50\%$ accuracy due to random guesses. In contrast, Flock succeeds on all considered instances, displaying that, while remaining invariant in probability, it can differentiate between non-isomorphic links, even with isomorphic relations.
@@ 5.2 Entity and relation prediction over knowledge graphs
Setup. We follow the protocol of Galkin et al. (2024); Zhang et al. (2024) and pretrain Flock on FB15k-237 (Toutanova & Chen, 2015), WN18RR (Dettmers et al., 2018), and CoDEx Medium (Safavi & Koutra, 2020). We then evaluate its zero-shot and finetuned inference performance with the test set of 54 KG datasets (see Table 15 for details). These KG datasets are extracted from diverse domains across three settings: inductive on nodes and relations (Inductive $e,r$), inductive on nodes (Inductive $e$), and transductive. Note that these settings differ only during finetuning setup; in zero-shot setup, all entities and relations are unseen. We choose two state-of-the-art KGFMs Ultra (Galkin et al., 2024) and Trix (Zhang et al., 2024) as baselines, since they are pretrained on the same pretraining graph mix to ensure a fair comparison. Following conventions in the literature (Zhu et al., 2021; Huang et al., 2023), for each triple $(h,r,t)$, we add the corresponding inverse triple $(t,r^{-1},h)$, where $r^{-1}$ is a fresh relation symbol. For entity prediction, we report both head and tail results for each triple on all datasets, except for three from Lv et al. (2020) where only tail results are available, following Zhang et al. (2024). For evaluation, we use the filtered ranking protocol (Bordes et al., 2013), reporting Mean Reciprocal Rank (MRR) and Hits@10 for entity prediction, and Hits@1 for relation prediction, as some datasets have fewer than $10$ relations. Detailed per-dataset results are shown in Appendix G.
Entity prediction results. We present the average entity prediction results in Table 2 (Q2). In the zero-shot setting, Flock consistently outperforms Ultra and Trix on all metrics, demonstrating its strong generalization on KGs over diverse domains. Notably, on Metafam (Zhou et al., 2023), a dataset designed to challenge models with conflicting and compositional relational patterns, Flock roughly doubles MRR over Ultra and achieves about a 40% MRR gain over Trix in zero-shot performance. These gains align with our hypothesis that probabilistic node–relation equivariance improves expressivity without sacrificing generalization. In the finetuning setting, we observe a similar pattern: Flock maintains a consistent improvement over all datasets except transductive splits, where the underlying KGs are generally larger. We hypothesize that this gap stems from random walk coverages. Unlike Ultra and Trix whose message passing guarantees a full $k$-hop neighborhood coverage over the queried node, Flock relies on sampling random walks, which may not fully cover the target nodes of interest.
@@
Relation prediction results. Table 3 presents the average MRR and Hits@1 for relation prediction results (Q2). We observe that Flock substantially outperforms all existing KGFMs across all categories in the zero-shot setting, achieving an $11.2\%$ relative improvement in MRR compared to the best baseline Trix. Flock shows a further performance boost of $12.8\%$ in the finetuned setting. We attribute this huge performance gain to Flock’s joint encoding of entities and relations during the updating step using the sequence encoder, while existing KGFMs, Ultra and Trix, require separate updating steps for entities and relations. This joint updating mechanism yields more holistic representations of both entities and relations with minimal information loss.
5.3 Scaling behavior of Flock
Pretraining graph mixing scaling. To assess whether Flock benefits from more pretraining graph and data (Q3), we follow the setup of Galkin et al. (2024), and pretrain Flock on an increasing number of KGs. We then evaluate them on all $41$ inductive benchmarks for a fair comparison. We present the detailed pretraining graph mix in Table 19. As shown in Figure 4(a), Flock’s generalization improves consistently as the number of pretraining KGs increases, exhibiting clear scaling behavior, which is a core characteristic of being a foundation model.
@@ 5.3 Scaling behavior of Flock
Number of ensembled predictions. To assess how test-time ensemble size $P$ affects performance (Q3), we take the pretrained Flock and run zero-shot entity prediction on 41 inductive KGs by increasing the number of ensembled passes. As shown in Figure 4(b), performance improves from 1 to 8 passes and then begins to plateau beyond 12. This indicates a clear scaling behavior: larger ensembles provide a more accurate estimate of the underlying node and relation distributions.
5.4 Impact of Test-Time Adaptation of Walk Counts
Table 4: Zero-shot entity prediction of Flock with and without adaptive test-time walks. We show the average number of entities $|V|$, triples $|E|$, base walks $n$, MRR, and Hits@10.
Dataset split Statistics Flock Flock w/o Adap.
@@ 5.4 Impact of Test-Time Adaptation of Walk Counts
Recall that we employ test-time adaptation of walk counts, which adaptively selects the base walk count $n$ based on the graph size, computed via the harmonic-mean rule shown in Equation 4 during inference. To answer Q4, we conduct an ablation study over this adaptive mechanism. Table 4 compares this adaptive setting with a fixed setting that uses $128$ base walks per sample for all datasets, matching the pretraining setup ($n_{\text{train}}=128$). As expected, the average selected base count $n$ is smaller on both inductive splits and larger on the transductive split, yet the adaptive mechanism improves performance across all settings. This is consistent with the intuition that adaptive $n$ scales up walks on larger KGs to improve coverage while allocating fewer walks on smaller KGs to reduce redundant visits; Flock maintains comparable visiting rates and coverage to those seen during pretraining, thereby producing representations closer to the pretraining distribution and resulting in consistent performance gains.
@@ 6 Conclusions
We introduced Flock, as a knowledge graph foundation model that respects probabilistic node-relation equivariance. Flock iteratively samples query-conditioned random walks from the input KG, records encountered nodes and relations via a recording protocol, and relies on a sequence processor and consensus protocol to obtain node and relation representations. We empirically evaluate Flock over 54 KGs across different domains for both entity and relation prediction tasks, demonstrating its superior zero-shot and finetuned performances. We further construct a synthetic dataset petals to validate our theoretical findings. One limitation is scalability (discussed in Appendix E): ensuring coverage of the sampled random walk in a large KG requires an extensive number of longer walks, which can quickly become computationally infeasible. A future direction is to develop approximation strategies (Chamberlain et al., 2023) that reduce the cost of random walk sampling while retaining Flock’s downstream performance.
@@ Acknowledgment
This work is supported in part by the National Research Foundation of Korea grant (No.RS-2024-00351212) and “HPC support” project funded by the Korea government (MSIT and NIPA). MB is supported by EPSRC Turing AI World-Leading Research Fellowship No. EP/X040062/1 and EPSRC AI Hub on Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI No. EP/Y028872/1.
Appendix A Details of the Petals benchmark
Motivated by this limitation, we introduce the Petals benchmark. Petals comprises 220 graphs, each paired with a link prediction query $(h,r,?)$ and a target set $\{t_{1},t_{2}\}$. While $t_{1}$ and $t_{2}$ are non-isomorphic, KGFMs enforcing relational invariance are unable to distinguish them, producing identical predictions. We empirically validate this property by evaluating the classification accuracy of marking $t_{1}$ as true and $t_{2}$ as false, reported in Table 1.
A.1 Structure of the studied KGs
Knowledge graphs in Petals follow a flower-like structure, parametrized by the number $c$ of ‘petals’, their length $l$ and the length $t$ of the ‘stem’ (see Figure 5 for visualization).
Edges. The nodes of the stem are connected in a consecutive manner by the same relation type $r_{0}$. Precisely, for each $i\in{1,\cdots,t}$, there exists an edge $(b_{i-1},r_{0},b_{i})$. Each petal $A^{(i)}$ is associated with two edge types $r^{(i)}_{1},r^{(i)}_{2}$, and is connected to the central node $b_{0}$ with links $\left(b_{0},r^{(i)}_{1},a_{1}^{(i)}\right)$ and $\left(b_{0},r^{(i)}_{2},a_{2}^{(i)}\right)$. The rest of the petal is connected with edges of type $r_{1}^{(i)}$ only, going from $a^{(i)}_{2j-1}$ to $a^{(i)}_{2j+1}$, and from $a^{(i)}_{2j}$ to $a^{(i)}_{2j+2}$. Finally, there are also edges linking $a^{(i)}_{2l-1}$ and $a^{(i)}_{2l}$ to $a^{(i)}_{2l+1}$. Therefore, the full set of edges can be characterized as:
$\cup\left(\bigcup_{i=1}^{c}\bigcup_{j=1}^{l-1}\left\{\!\left(a^{(i)}_{2j-1},r^{(i)}_{1},a^{(i)}_{2j+1}\right),\!\left(a^{(i)}_{2j},r^{(i)}_{1},a^{(i)}_{2j+2}\right)\right\}\right)$
$\cup\left(\bigcup_{i=1}^{c}\left\{\!\left(a^{(i)}_{2l-1},r^{(i)}_{1},a^{(i)}_{2l+1}\right),\!\left(a^{(i)}_{2l},r^{(i)}_{1},a^{(i)}_{2l+1}\right)\right\}\right)$
We select each of the types $r^{(i)}_{1}$ and $r^{(i)}_{2}$ from the set of considered relations $R=\{r_{1},\dots,r_{|R|}\}$ so that any relation-invariant model will equate all petals (i.e. so that for each pair of petals, there is an automorphism taking one to another). For instance, Figure 5 displays a cyclic pattern, in which $r^{(i)}_{2}=r^{(i+1)}_{1}$. Such symmetry causes all petals to be isomorphic, and leads to the inability of KGFMs to distinguish between the relations inside them.
Link prediction instances. Although the petals are isomorphic to each other, given the asymmetry of edge types from $b_{0}$ to $a^{(i)}_{1}$ and $a^{(i)}_{2}$, the nodes within a single petal generally can be distinguished. Therefore, for each graph with the structure as described above, we randomly sample one of the stem nodes $b_{s}$, and ask the link prediction query $(b_{s},r_{0},?)$. For the target nodes, we randomly select petal index $i$ and distance $j$ from the central node $b_{0}$, and consider the predictions for $a^{(i)}_{2j-1}$ and $a^{(i)}_{2j}$. For example, Figure 5 shows the case when $b_{s}=b_{0}$, $i=1$ and $j=1$, where the query is $(b_{0},r_{0},?)$ and we are interested in the scores for $a^{(1)}_{1}$ and $a^{(1)}_{2}$.
A.2 Parameters and generation
Appendix B Methodology - details

### v2

@@ 4.1 Flock
4.2 Theoretical analysis
Expressivity. Following the notion of probabilistic expressivity introduced by Abboud et al. (2021), we say that a Flock model $X_{\theta}$ is a universal approximator of link invariant functions over $\mathbb{K}_{n,m}$ if for any link invariant $\varphi:\mathbb{K}_{n,m}\rightarrow(V\times R\times V\rightarrow[0,1])$ and any $\epsilon,\delta>0$, there exists a choice of the network parameters $\theta$ and the length of the sampled random walks $\ell$, such that:
@@ 4.2 Theoretical analysis
All proofs are in Appendix˜B. To offer an intuition behind the result, we provide a proof sketch.
Proof sketch. A sufficiently long random walk will cover all edges of the graph with high probability. Then, from its anonymized version, assigning unique positional identifiers to every node and relation, one can reconstruct the input graph, up to isomorphism. Thus, with a sufficiently expressive sequence processor, Flock can approximate any link-invariant function.
@@
Proof sketch. Since each of these components is invariant (in probability), and invariance of individual component is preserved under composition, we have that Flock is invariant.
@@
Any Flock model with components as outlined in this section, and detailed in Appendix˜A is invariant in probability.
@@ 5 Experiments
We test Flock over a wide range of KGs for both entity and relation predictions, aiming to answer:
@@
    How does performance scale with the sizes of pretraining graph mix and test-time ensemble?
@@
    What is the impact of choices of each component on the behavior and performance of Flock?
We further provide detailed scalability analysis in Appendix˜D and Appendix˜E, and comparisons against current KGFMs augmented with noise injection in Appendix˜F. Experiment details and hyperparameters are in Appendix˜I.
5.1 Synthetic dataset
@@
Figure 3: A KG in Petals. KGFMs with relational invariants must equate $r_{1}$ and $r_{2}$, thus predicting the same scores for dashed queries with $r_{0}$.
Setup.
To validate the limitations of KGFMs with node-relation equivariance (Q1), we construct a synthetic benchmark Petals. It contains $220$ instances, each including: (1) a KG $G=(V,E,R)$ consisting of a ‘central’ node $s$, a ‘stem’ $T\subset V$ with query relation $r_{0}$, and multiple cyclic ‘petals’, each ‘colored’ with a different pair of relations in $R\setminus\{r_{0}\}$, (2) an entity prediction query $(h,r_{0},?)$ with $h\in\{s\}\cup T$, and (3) two candidate targets $t_{1}$ and $t_{2}$ from the same ‘petal’, located at the same distance from $s$. An example is in Figure˜3. See Appendix˜C for more details.
Petals is designed such that each instance always admits non-trivial automorphisms, meaning that swapping relations occurring in the same ‘petal’ results in an isomorphic KG. Consequently, any model computing relation invariants will not be able to distinguish between potential links $(s,r_{0},t_{1})$ and $(s,r_{0},t_{2})$. However, the samples are constructed so that these links are not isomorphic from the graph perspective, making them distinguishable for general link-invariant functions. We say a model solves an instance if it can classify $(s,r_{0},t_{1})$ as true and $(s,r_{0},t_{2})$ as false. For empirical validation, we train Ultra (Galkin et al., 2024), $\textsc{Motif}({\mathcal{F}}_{\text{Path}}^{3})$ (Huang et al., 2025), Trix (Zhang et al., 2024), and Flock from scratch and validate them on the training instances.
Table 1: Petals accuracies.
@@
Results.
The results are provided in Table˜1. As expected, all existing KGFMs relying on learning deterministic relational invariants fail to distinguish between the candidate target triplets completely, achieving $50\%$ accuracy due to random guesses. In contrast, Flock succeeds on all considered instances, displaying that, while remaining invariant in probability, it can differentiate between non-isomorphic links, even with isomorphic relations.
@@ 5.2 Entity and relation prediction over knowledge graphs
Setup.
To answer Q2, we follow the protocol of Galkin et al. (2024); Zhang et al. (2024) and pretrain Flock on FB15k-237 (Toutanova & Chen, 2015), WN18RR (Dettmers et al., 2018), and CoDEx Medium (Safavi & Koutra, 2020). We then evaluate its zero-shot and finetuned inference performance with the test set of 54 KGs (see Appendix˜I for details). These KGs are extracted from diverse domains across three settings: inductive on nodes and relations (Inductive $e,r$), inductive on nodes (Inductive $e$), and transductive. Note that these settings differ only during finetuning setup; in zero-shot setup, all entities and relations are unseen. We choose state-of-the-art KGFMs Ultra (Galkin et al., 2024) and Trix (Zhang et al., 2024) as baselines, as they are pretrained on the same KGs, to ensure a fair comparison. For evaluation, we use the filtered ranking protocol (Bordes et al., 2013), reporting mean reciprocal rank (MRR) and Hits@10 for entity prediction, and Hits@1 for relation prediction, as some KGs have fewer than 10 relations. Per-dataset results are in Appendix˜I.
Entity prediction.
Table˜2 shows the entity prediction results. In the zero-shot setting, Flock consistently outperforms Ultra and Trix, demonstrating strong generalization across diverse domains. Notably, on Metafam (Zhou et al., 2023), a dataset designed to challenge models with conflicting and compositional relational patterns, Flock roughly doubles MRR over Ultra and achieves $\approx$ 40% gain in MRR over Trix in the zero-shot setting. We find that Flock distinguishes structurally similar but semantically conflicting relations while Ultra fails (Appendix˜G), explaining the gain. These findings align with our hypothesis that probabilistic equivariance improves expressivity without sacrificing generalization. In the finetuning setting, we observe a similar pattern: Flock maintains a consistent improvement over all datasets except transductive ones, where KGs are generally larger. We hypothesize that the gap stems from walk coverages. Unlike Ultra and Trix, whose message passing guarantees a full neighborhood coverage around the queried node, Flock relies on sampled walks that may not fully cover target nodes of interest. We find that Flock favors sparse KGs (Appendix˜H), consistent with this hypothesis as random walks cover sparse graphs faster.
Relation prediction.
Table˜3 presents the relation prediction results. Flock substantially outperforms all existing KGFMs across all categories in the zero-shot setting, achieving an $11.2\%$ relative improvement in MRR compared to the best baseline Trix. Flock shows a further performance boost of $12.8\%$ in the finetuned setting. We attribute this huge gain to Flock’s joint encoding of entities and relations during the updating step via the sequence encoder, while existing KGFMs, Ultra and Trix, have separate update steps for entities and relations. This joint update mechanism yields more holistic representations of both entities and relations with minimal information loss.
5.3 Scaling analysis
Size of pretraining graph mix.
To assess whether Flock benefits from more pretraining graph and data (Q3), we follow the setup of Galkin et al. (2024), and pretrain Flock on an increasing number of KGs. We then evaluate them on all $41$ inductive benchmarks for a fair comparison. We present the detailed pretraining graph mix in Table˜16. As shown in Figure˜4(a), Flock’s generalization improves consistently as the number of pretraining KGs increases, exhibiting clear scaling behavior, which is a core characteristic of being a foundation model.
@@ 5.3 Scaling analysis
Number of ensembled predictions. To assess how test-time ensemble size $P$ affects performance (Q3), we take the pretrained Flock and run zero-shot entity prediction on 41 inductive KGs by increasing the number of ensembled passes. As shown in Figure˜4(b), performance improves from 1 to 8 passes and then begins to plateau beyond 12. This indicates a clear scaling behavior: larger ensembles provide a more accurate estimate of the underlying node and relation distributions.
5.4 Ablation studies
Table 4: Ablation study of adaptive test-time walks with zero-shot entity prediction task. We show the average number of entities $|V|$, triples $|E|$, base walks $n$, MRR, and Hits@10.
Dataset split Statistics Flock Flock w/o Adap. Equation˜1
@@ 5.4 Ablation studies
Table 5: Detailed ablation study with zero-shot entity prediction task. For the transductive split, considering resource limits, we test NELL995, NELL23k, WDsinger, ConceptNet100k, and YAGO310.
Inductive $e,r$ Inductive $e$ Transductive Total Avg Model (23 graphs) (18 graphs) (5 graphs) (46 graphs) MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 Flock ($\ell=128$) 0.369 0.554 0.456 0.604 0.360 0.542 0.395 0.567 w/o non-backtracking 0.370 0.549 0.456 0.605 0.334 0.499 0.386 0.551 $\ell=64$ 0.372 0.556 0.459 0.606 0.351 0.534 0.394 0.565 $\ell=256$ 0.360 0.548 0.458 0.605 0.338 0.508 0.385 0.553 w/o diverse starts 0.360 0.539 0.448 0.596 0.319 0.488 0.385 0.553 transformer $f_{\theta}$ 0.356 0.542 0.410 0.591 0.312 0.477 0.359 0.537 w/o weighted consensus 0.351 0.526 0.448 0.593 0.361 0.515 0.387 0.545
Setup.
To understand the impact of each design choice on the performance and behavior of Flock (Q4), we conduct a series of ablation studies spanning random walks, sequence processor and the consensus protocol, in the entity prediction task in the zero-shot setting.
Adaptive test-time walks. Recall that we employ test-time adaptation of walk counts, which adaptively selects the base walk count $n$ based on the graph size, computed via the harmonic-mean rule shown in Equation˜1 during inference. Table˜4 compares this adaptive setting with a fixed setting that uses $128$ base walks per sample for all datasets, matching the pretraining setup ($n_{\text{train}}=128$). As expected, the average selected base count $n$ is smaller on both inductive splits and larger on the transductive split, yet the adaptive mechanism improves performance across all settings. This is consistent with the intuition that adaptive $n$ scales up walks on larger KGs to improve coverage while allocating fewer walks on smaller KGs to reduce redundant visits; Flock maintains comparable visiting rates and coverage to those seen during pretraining, thereby producing representations closer to the pretraining distribution and resulting in consistent performance gains.
Non-backtracking walks. Flock employs non-backtracking uniform random walk, which has an effect of faster exploration and coverage of distant regions (Alon et al., 2007). In Table˜5, we compare this with uniform walk that may backtrack and hence is slower in global exploration. While non-backtracking does not alter results much on inductive splits, it significantly improves performance on the transductive split. This is consistent with the idea that improving coverage especially benefits performances on large KGs, which Flock achieves via non-backtracking.
Walk lengths. Flock uses random walks of length $\ell=128$, a choice made by finding the lowest $\ell$ reliably visiting target node and relation on various KGs. Table˜5 compares this with shorter and longer walks by a factor of two. As expected from coverage, shorter walks show degraded results on the transductive split with large KGs. Longer walks are overall worse, which is explained by higher learning complexity of the sequence processor that has a small hidden dimension (64) for scalability. Flock finds a balance of coverage and learnability, achieving robust results on diverse KGs.
Diverse starting locations of walks. We recall that Flock uses a diversification strategy of starting locations of walks, with $n$ walks from the query node, $n$ walks from random relation, and $n$ walks from random node, adding up to $3n$ walks capturing both local context near query and global information of KG. Table˜5 compares this against all $3n$ walks starting at the query node. As expected, this causes degradations on all splits, showing the benefit of using both local and global information.
Sequence processor. Flock uses a sequence processor $f_{\theta}$ with bidirectional GRU. In Table˜5, we compare this against a transformer $f_{\theta}$ with a similar SwiGLU-RMSNorm architecture (Dubey et al., 2024) and parameter count. This alternative does not deliver good results, which is explained by the restrictions on model scales that are enforced to scale to large KGs. Flock benefits from reasoning efficiency of GRU in limited parameter regime, gaining good performance and scalability together.
Consensus protocol. Flock uses softmax-weighted averaging to pool sequence processor outputs into state updates for nodes and relations, under the intuition that this can suppress uninformative proposals from the sequence processor better than simple, unweighted averaging. Table˜5 provides a comparison, showing that weighted consensus outperforms the unweighted counterpart. This verifies our intuition on how the design of the consensus protocol strengthens Flock.
@@ 6 Conclusions
We introduced Flock, a knowledge graph foundation model that respects probabilistic node-relation equivariance. Flock iteratively samples query-conditioned random walks, records encountered nodes and relations via a recording protocol, and relies on a sequence processor and consensus protocol to obtain node and relation representations. We evaluate Flock on 54 KGs across different domains for both entity and relation prediction, demonstrating superior zero-shot and finetuned performance. We further construct a synthetic dataset petals to validate our theoretical findings. One limitation is scalability (Appendix˜E): ensuring coverage of the sampled random walk in large KGs requires many long walks, which can quickly become computationally infeasible. A future direction is to develop approximation strategies (Chamberlain et al., 2023; Łącki et al., 2020) that reduce the cost of random walk sampling while retaining Flock’s downstream performance. Another avenue for future work is studying the families of approximable functions when the walk lengths are restricted, for example based on connections to subgraph-based reconstructions (Cotta et al., 2021).
@@ Acknowledgment
This work was in part supported by the National Research Foundation of Korea (RS-2024-00351212 and RS-2024-00436165) and the Institute of Information & Communications Technology Planning & Evaluation (IITP) (RS-2024-00509279, RS-2022-II220926, RS-2022-II220959, and RS-2019-II190075) funded by the Korean government (MSIT). Computational resources were in part provided by the “HPC support” project funded by MSIT and NIPA. MB is supported by EPSRC Turing AI World-Leading Research Fellowship No. EP/X040062/1 and EPSRC AI Hub on Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI No. EP/Y028872/1.
Ethics statement
This work introduces a probabilistic framework for knowledge graph foundation models, aiming to improve the generalization in zero-shot link prediction. Our contributions are methodological and theoretical, with evaluations performed on publicly available benchmarks and a synthetic dataset designed to validate our theoretical results. We do not anticipate any direct ethical risks associated with this approach. We acknowledge and adhere to the ICLR Code of Ethics.
Reproducibility statement
We make every effort to ensure the reproducibility of the experiments in our paper. We release a codebase at https://github.com/jw9730/flock with training and evaluation scripts for Flock, including pretraining scripts and checkpoints on FB15k-237, WN18RR, and CoDEx Medium, evaluation over 54 KGs, and the synthetic dataset Petals generator. All architectural details needed to reimplement the method, including the random-walk sampler, recording protocol, sequence processor, and consensus protocol, are specified in Appendix˜A, and our theoretical claims are supported with complete proofs in Appendix˜B. We additionally include further experimental details in Appendix˜I.
Appendix A Methodology - details